If the sides of a square are each increased by 4 inches, the area becomes
121 square inches. Find the perimeter of the original square.

Answer by mananth(16946)

If the sides of a square are each increased by 4 inches, the area becomes
121 square inches. Find the perimeter of the original square.
..
let side =x
x+4 is the increased length
...
Area = L^@
(x+4)^2=121
x^2+8x+16=121
x^2+8x-105=0
x^2+15x-7x-105=0
x(x+15)-7(x+15)=0
(x+15)(x-7)=0
x= 7
perimeter = 4L = 28 inches
...
m.ananth@hotmail.ca